If a team won 72percent72percent of their games. What is the ratio of games won to number of games lost?
Found 2 solutions by richwmiller, dfrazzetto:
Answer by richwmiller(17219)

72/28=36/14=18/7

Answer by dfrazzetto(283)


Did you mean to write 72 percent of 72 percent of their games?
.72(.72) = 51.84% won
100- 51.84 = 48.16% lost
ratio = 51.84/48.16